OpenStudy (anonymous):

PLEASE!!!!!!!!!!!! how do you find the definite integral from 4 to 5 of x(x^2 + 3y)^(1/2) dy ? ???

zepdrix (zepdrix):

\[\large \large \int\limits_4^5 x \sqrt{x^2+3y}dy\]Hmm, maybe just apply a U-substitution so you can figure out what's going on here :) Remember that we're treating X as a constant since we're integrating with respect to y. It will even be constant when we make a substitution. \[\large \color{orangered}{u=x^2+3y}\]

zepdrix (zepdrix):

\[\large du=0+3 dy\]Understand how that du came about? :D

zepdrix (zepdrix):

Dividing both sides by 3 gives us, \[\large \color{orangered}{\frac{1}{3}du=dy}\] If we plug the orange pieces into our integral we get,\[\large \int\limits_{y=4}^5 x \sqrt u \left(\frac{1}{3}du\right)\] Pulling the constants outside gives us,\[\large \frac{1}{3}x \int\limits_{y=4}^5 u^{1/2}du\]

OpenStudy (anonymous):

omg how did you know to set u to that instead of x?? i started to use integration by parts way and its not working T_T

zepdrix (zepdrix):

D: See how the differential at the end is a \(dy\)? So we don't really care much about X in this integral at all since there is no \(dx\) present. So we treat all X's as constants.

zepdrix (zepdrix):

If it helps, think of the integral like this,\[\large \int\limits_4^5 a \sqrt{a^2+3y}\;dy\]Where \(a\) is just some constant that we don't have to worry about while integrating. :o

OpenStudy (anonymous):

ohh so we can just take x out of the integral before we even start the problem? so the integrand would only be radical x^2 + 3y? is there no use for integration nby parts now lol sorry T.T

zepdrix (zepdrix):

Yah we don't need to do any by-parts c: That sounds messy anyway :U hah!

OpenStudy (anonymous):

ok thank you so much!!!!!! !!!!!